OpenStudy (anonymous):

One number is 4 more than another, and their sum is 60. What is the smaller number? If x = the larger number and y = the smaller number, then which of the following systems could be used to solve the problem?

OpenStudy (anonymous):

x + y = 60 and -x + y = 4 x + y = 60 and x - y = 4 x + y = 60 and x + y = 4

OpenStudy (anonymous):

which of the equations describes x is larger than y by 4?

OpenStudy (ajprincess):

x=y+4-(larger number is 4 more than smaller number) x+y=60

OpenStudy (anonymous):

The second one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks!